### v1

@@ 3.2 Construction of a Separable Approximation
Assume that $V$ possesses the $\gamma$-decaying sensitivity property with $\gamma(l)=C\rho^{l}$ for some $C>0$ and $\rho\in(0,1)$, and that there exists a growth bound $r\colon\mathbb{N}\to\mathbb{N}$ for the underlying graph, satisfying $r(k)\leq\hat{C}\mu^{k}$, $k\in\mathbb{N}$, for some $\hat{C}>0$ and $\mu\in[1,\rho^{-1})$. Then for all $l\leq s$ and all $x\in\Omega$ it holds
@@
Let $l\leq s$, $D_{l}$ as in (6), $i\in\{1,\dots,s\}$, and $\delta\in(\rho\mu,1)$. Then it holds
@@
$\lVert D_{l}\rVert_{\infty}=\max_{i\in[s]}\sum_{j\in[s]}\lvert D_{l}[i,j]\rvert\leq C\sup_{k\in\mathbb{N}}\left\{r(k)\left(\frac{\rho}{\delta}\right)^{k}\right\}\frac{\delta}{1-\delta}\delta^{l}$
$\leq C\hat{C}\sup_{k\in\mathbb{N}}\left\{\mu^{k}\left(\frac{\rho}{\delta}\right)^{k}\right\}\frac{\delta}{1-\delta}\delta^{l+1}\leq C\hat{C}\frac{\delta}{1-\delta}\delta^{l+1}=:\widetilde{C}\delta^{l+1}.$
@@
$\lvert V(x)-\Psi_{l}(x)-V(0)\rvert\leq\lVert x\rVert_{2}^{2}\lVert D_{l}\rVert_{2}\leq\sqrt{\lVert A\rVert_{1}\lVert A\rVert_{\infty}}\leq\widetilde{C}\delta^{l+1}.$
@@
Assume that $V$ possesses the $\gamma$-decaying sensitivity property with $\gamma(l)=C(l+1)^{-\alpha}$ for some $C>0$ and $\alpha>1$. Furthermore, assume that there exists a growth bound $r\colon\mathbb{N}\to\mathbb{N}$ for the underlying graph and some $\beta<\alpha-1$ such that $r\in\mathcal{O}(l^{\beta})$ for $l\to\infty$. Then for all $l\leq s$ it holds that for all $x\in\Omega$
@@
Theorems 3.4 and 3.5 are stated for a fixed dimension $n$ and a fixed number of subsystems $s$. However, the resulting approximations are dimension-independent, as the approximation quality does not degrade with increasing $s$ or $n$. Consequently, if we consider a family of problems with increasing dimension and the assumptions are satisfied uniformly with respect to the dimension, the approximations obtained are also uniform in the dimension.
@@
$f_{\varnothing,c}=f(c),\quad f_{u,c}(x_{u})=f((x;c))-\sum_{v\subsetneq u}f_{v,c}(x_{v}).$ (10)
The key idea behind anchored decomposition is to express an high-dimensional function $f(z_{1},\ldots,z_{s})$ as a sum of terms that depend on a progressively larger number of variables, starting from simpler, lower-dimensional components. This decomposition allows simplifying the computation by focusing on low-dimensional interactions (e.g., pairwise or three-variable interactions) when possible, making high-dimensional problems more tractable. The $d$-separable decomposition shares some terms of the anchored decomposition and neglects those which include indices far away according to the graph distance.
@@
In summary, anchored decomposition and separable decomposition share a common goal of simplifying high-dimensional functions by focusing on lower-dimensional interactions. While anchored decomposition systematically includes all terms based on variable subsets, separable decomposition selectively neglects terms involving distant variables, leveraging graph-based distances. This distinction allows separable decomposition to achieve greater computational efficiency by emphasizing localized interactions, making it particularly suited for functions with decaying sensitivity over distance.
@@ 4 Sensitivity decay in linear-quadratic optimal control
Based on 2.1, we define the state $x\in\mathbb{R}^{n}$ as being composed of $s\leq n$ substates $x=(z_{1},\dots,z_{s})$, where each substate $z_{j}$ may have an associated control $u_{j}\in\mathbb{R}^{m_{j}}$ for $j\in\{1,\dots,s\}$. Whenever a substate does not have an associated control, we set $m_{j}=0$. The collection of all controls is denoted by $u=(u_{1},\dots,u_{s})\in\mathbb{R}^{m}$, with $m=\sum_{j=1}^{s}m_{j}$. Given matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $Q\in\mathbb{R}^{n\times n}$, $Q\succeq 0_{n}$, and $R\in\mathbb{R}^{m\times m}$, $R\succ 0_{m}$, we consider an infinite-horizon optimal control problem formulated either in discrete time
$\begin{split}\min\;&J(x_{0},u)=\sum_{k=0}^{\infty}\ell(x(k),u(k)),\quad\text{s.t.}\;x(k+1)=f(x(k),u(k)),\;x(0)=x_{0},\end{split}$ (12)
@@
[…]
@@
where $\mathcal{U}\coloneqq L^{\infty}(\mathbb{R}_{\geq 0},\mathbb{R}^{m})$. In the LQR framework, it is well-known that the optimal value function $V$ is quadratic, taking the form $V(x)=x^{\top}Px$, where $P$ satisfies the Discrete Algebraic Riccati Equation for discrete-time systems, and the Continuous Algebraic Riccati Equation for continuous-time systems. As described in Setting 2.1, the graph $\mathcal{G}$ associated with such a problem is determined by the subsystems and their mutual influences. For the set of vertices $\{1,\dots,s\}$, the graph $\mathcal{G}$ has an edge $(j,i)$ if
@@
As highlighted in Corollary 3.1, the decaying sensitivity of the value function is associated with the SED property of the Riccati solution. In the first subsection, we address the discrete-time setting, where we revisit results previously derived in [29]. In the second subsection, we transition to the continuous-time setting and, under appropriate assumptions on the problem data, demonstrate an off-diagonal decay property of the Riccati solution. We introduce some notions that will be used in the following.
@@
The pair $(A,B)$ is called stabilizable if there exists a feedback matrix $K\in\mathbb{R}^{n\times m}$ such that $A-BK$ is stable.
@@ 4.1 Discrete-Time Linear Quadratic Regulator
    $(A,B)$ is uniformly stabilizable, i.e., Definition 4.1 holds with $\|K\|\leq L$ and $C=A-BK$ satisfies (15),
@@
$|K[i,j]|\leq C_{K}\rho_{K}^{\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)},$
@@
Note that Proposition 4.1 is stated for the case $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=|i-j|$, but its proof extends naturally to any graph where $|\mathcal{G}_{i,1}|=|\{j\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq 1\}|$ remains bounded independently of the dimension for all $i\in\{1,\dots,s\}$. For example, if $B=Q=R=I_{n}$ and $A$ is a $p$-banded matrix for some $p\in\mathbb{N}$, it follows that $|\mathcal{G}_{i,1}|\leq 2p+1$ for any dimension and any node $i$.
@@ 4.2 Continuous-Time Linear Quadratic Regulator
First, observe that $A^{2}+\frac{c}{\gamma}I_{n}>0$, ensuring that the matrix square root in the expression for $P$ is well defined. Furthermore, since $\sqrt{\lambda^{2}+\frac{c}{\gamma}}>|\lambda|$ and $\gamma>0$, the matrix in (18) is positive definite. Substituting expression (18) into the ARE (17) confirms that the expression for $P$ satisfies the equation. ∎
The expression (18) can be reformulated as $P=f(A)$, where
@@
remains banded with the same bandwidth, and its spectrum is contained in $[-1,1]$. Moreover, $f(A)=f(\psi^{-1}(\widetilde{A}))$. Accordingly, we focus on bounding the off-diagonal entries of the matrix function $F(\widetilde{A})$ where for $x\in\mathbb{C}$
@@
and $\widetilde{A}$ is banded with spectrum in $[-1,1]$. The following useful result is adapted from [4, Theorem 2.2].
@@
Let $g:\mathbb{C}\rightarrow\mathbb{C}$ be a function analytical in the interior of an ellipse $\mathcal{E}_{\chi}$ with foci at ±$\pm 1$ and sum of the half axes $\chi>1$, and continuous on $\mathcal{E}_{\chi}$. Additionally, let $A\in\mathbb{C}^{n\times n}$ be symmetric, $p$-banded, and such that $[-1,1]$ contains its spectrum, and let $\rho:=\chi^{-\frac{1}{p}}\in(0,1)$. Then, for all pairs of indices $(i,j)$ such that $i\neq j$, we have that
@@
and $\mathcal{E}_{\chi}$ is the ellipse with foci at ±$\pm 1$ and sum of the half axes $\chi>1$.
@@
We have $f(A)=f(\psi^{-1}(\widetilde{A}))=F(\widetilde{A})$ and the function $F(x)$ (given in (20)) satisfies the conditions of Theorem 4.2. In particular, it is analytical in any ellipse with foci in $-1$ and $1$ that do not include, in its interior part, the points
@@
Thus, we consider the ellipse $\mathcal{E}_{\chi}$ with foci in $-1$ and $1$ passing through the points $x_{1/2}$. The sum of the half axes of $\mathcal{E}_{\chi}$ is given by
@@ 5 Efficient Neural Network Representation for Separable Structures
In this section, we examine how neural networks can efficiently represent functions exhibiting decaying sensitivity. While neural networks are universal approximators, they typically require an exponential growth in neurons and parameters as input dimensionality increases. However, we demonstrate that for functions $V$ possessing a separable approximation, this curse of dimensionality can be mitigated. Under appropriate conditions, such functions can be represented with neural networks requiring a number of parameters which only grows polynomially. This result is consistent with existing literature showing that certain structural properties enable polynomial parameter scaling with dimension. For example, a similar result holds for compositional functions (cf. [13]) and in particular, for the sub-class of separable functions. The approximation of separable functions with neural networks has been discussed in [10, 28] in the context of (control) Lyapunov functions. While the approach and the neural network construction in the following theorem is similar, there are two major differences compared to these works. The first one is that [10, 28] considers strictly separable functions, meaning that each component $x_{i}$ of the state $x$ appears in exactly one subsystem. This is not the case in the framework of this paper, since we achieve separability by considering overlapping graph neighborhoods, whence we cannot directly apply the results in [10, 28] towards our setting. On the other hand, in [10, 28] the authors used a linear transformation as first hidden layer for the computation of suitable substates to achieve separability. This is not necessary in our case, since for any given length $l$ of the graph-neighborhoods, the vectors $z_{j,l}$ are uniquely determined by the graph $\mathcal{G}$.
@@
respectively, where $C>0$ and $d>0$ are independent of $s$ and $n$, but determined by $\varepsilon$, $R$, the parameters $C$ and $\rho$ determining the exponential decay in Theorem 3.4, and the dimension of graph-neighborhoods in $\mathcal{G}$.
@@
Let $l\geq\ln\left(\frac{\epsilon}{\widetilde{C}}\right)\ln^{-1}(\delta)-1$, where $\widetilde{C}$ and $\delta$ are defined as in Theorem 3.4. Let $\Psi_{l}(x)$ be as defined in (4) and (5). Then, by Theorem 3.4, we obtain that for all $x\in\Omega$
@@
Thus, it is sufficient to show the existence of a neural network that approximates $\Psi_{l}$ on $\Omega$ with error bounded by $\epsilon$. We begin by constructing subnetworks for approximating $\Psi_{l}^{j}$, $1\leq j\leq s$. Recall that $\Psi_{l}^{j}$ depends on $z_{j,l}\in\mathbb{R}^{d_{j,l}}$. By Theorem 2.1 in [19] and its extension to cubes with length $R$ in [13], there exists a constant $\mu_{d_{j,l}}$ such that any neural network $W_{j}(\cdot;\theta_{j})\colon\mathbb{R}^{d_{j,l}}\to\mathbb{R}$ consisting of one hidden layer with $M\in\mathbb{N}$ neurons and activation function $\sigma$ satisfies
$\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\leq\mu_{d_{j},l}\,\tilde{R}\,{M}^{-\frac{1}{d_{j,l}}}\|\Psi_{l}^{j}\|_{C^{1}([-R,R]^{d_{j,l}})},$
where $\tilde{R}\coloneqq\max\{1,R\}$. Since, by assumption, $V$ has a $C^{1}$-norm bounded by $C_{V}$ on $[-R,R]^{n}$, it follows from (5) that $\Psi_{l}^{j}$ has $C^{1}$-norm bounded by $2C_{V}$ on $[-R,R]^{d_{j,l}}$, leading to
$\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\leq 2C_{V}\mu_{d_{j},l}\,\tilde{R}\,{M}^{-\frac{1}{d_{j,l}}}.$ (22)
@@
$\begin{split}&\inf_{\theta}|W(x;\theta)-\Psi_{l}(x)|_{\infty,\Omega}\leq\sum_{j=1}^{s}\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\\ \leq&2sC_{V}\mu_{d}\,R\,{M}^{-\frac{1}{d}},\end{split}$
where $d\coloneqq\max_{j}d_{j,l}$. Solving for $M$ yields $M\geq s^{d}(2C_{V}\mu_{d}R)^{d}\frac{1}{\varepsilon^{d}}$. To establish the existence of a parameter $\theta$ that achieves the desired error bound $\varepsilon$, we choose
$\bar{M}=s^{d}(4C_{V}\mu_{d}R)^{d}\frac{1}{\varepsilon^{d}}=:s^{d}C^{d}$
@@
Theorem 5.1 states that, under the assumption of exponentially decaying sensitivity, neural networks can approximate $V$ with a number of neurons and parameters that grows polynomially with respect to the number of subsystems $s$, with exponent $d$. The maximum dimension of a graph-neighborhood $d$ depends on the choice of $l$. Therefore, for a family of problems with increasing dimension, where, for any fixed $l$, the maximum neighborhood dimension $d$ remains constant, Theorem 5.1 describes a situation where the curse of dimensionality is avoided. This applies, for example, to a sequential graph with one-dimensional subsystems, where $d=l+1$ for any number of subsystems. However, it is important to note that Theorem 5.1 does not address the required growth of the network in relation to the accuracy $\varepsilon$.

### v2

@@ 3.2 Construction of a Separable Approximation
Assume that $V$ possesses the $\gamma$-decaying sensitivity property with $\gamma(l)=C\rho^{l}$ for some $C>0$ and $\rho\in(0,1)$, and that there exists a growth bound $r\colon\mathbb{N}\to\mathbb{N}$ for the underlying graph, satisfying $r(k)\leq\widehat{C}\mu^{k}$, $k\in\mathbb{N}$, for some $\widehat{C}>0$ and $\mu\in[1,\rho^{-1})$. Then for all $l\in\mathbb{N}$ and all $x\in\Omega$ it holds
@@
Let $l\in\mathbb{N}$, $D_{l}$ as in (6), $i\in\{1,\dots,s\}$, and $\delta\in(\rho\mu,1)$. Then it holds
@@
$\lVert D_{l}\rVert_{\infty}=\max_{i\in[s]}\sum_{j\in[s]}\lvert D_{l}[i,j]\rvert\leq C\sup_{k\in\mathbb{N}}\left\{r(k)\left(\frac{\rho}{\delta}\right)^{k}\right\}\frac{\delta^{l+1}}{1-\delta}$
$\leq C\widehat{C}\sup_{k\in\mathbb{N}}\left\{\mu^{k}\left(\frac{\rho}{\delta}\right)^{k}\right\}\frac{\delta^{l+1}}{1-\delta}\leq C\widehat{C}\frac{\delta^{l+1}}{1-\delta}\delta^{l+1}=:\widetilde{C}\delta^{l+1}.$
@@
$\lvert V(x)-\Psi_{l}(x)-V(0)\rvert\leq\lVert x\rVert_{2}^{2}\lVert D_{l}\rVert_{2}\leq\sqrt{\lVert D_{l}\rVert_{1}\lVert D_{l}\rVert_{\infty}}\leq\widetilde{C}\delta^{l+1}.$
@@
Assume that $V$ possesses the $\gamma$-decaying sensitivity property with $\gamma(l)=C(l+1)^{-\alpha}$ for some $C>0$ and $\alpha>1$. Furthermore, assume that there exists a growth bound $r\colon\mathbb{N}\to\mathbb{N}$ for the underlying graph and some $\beta<\alpha-1$ such that $r\in\mathcal{O}(l^{\beta})$ for $l\to\infty$. Then for all $l\in\mathbb{N}$ it holds that for all $x\in\Omega$
@@
Theorems 3.4 and 3.5 are stated for a fixed dimension $n$ and a fixed number of subsystems $s$. However, the resulting approximations are dimension-independent, as the approximation quality does not degrade with increasing $s$ or $n$. Consequently, if we consider a family of problems with increasing dimension and the assumptions are satisfied uniformly with respect to the dimension, i.e., each member of the family admits the same $\gamma$-decaying sensitivity and the same growth-bound $r$, then the resulting approximations are also uniform in the dimension.
@@
$f_{\varnothing,c}=f(c),\quad f_{u,c}(x_{u})=f((x_{u};c))-\sum_{v\subsetneq u}f_{v,c}(x_{v}).$ (10)
Here the notation $(x_{u};c)\in\mathbb{R}^{d}$ denotes the vector obtained by inserting the anchor $c_{i}$ into every coordinate $i\notin u$ and the free variable $x_{i}$ into every coordinate $i\in u$. Equivalently, if
$y_{i}\;=\;\begin{cases}x_{i},&i\in u,\\ c_{i},&i\notin u,\end{cases}\quad i=1,\dots,d,$
then
$(x_{u};c)\;=\;(y_{1},\dots,y_{d}),\qquad f\bigl{(}(x_{u};c)\bigr{)}\;=\;f(y_{1},\dots,y_{d}).$
The key idea behind anchored decomposition is to express a high-dimensional function $f(z_{1},\ldots,z_{s})$ as a sum of terms that depend on a progressively larger number of variables, starting from simpler, lower-dimensional components. This decomposition allows simplifying the computation by focusing on low-dimensional interactions (e.g., pairwise or three-variable interactions) when possible, making high-dimensional problems more tractable. The $d$-separable decomposition shares some terms of the anchored decomposition and neglects those which include indices far away according to the graph distance.
@@
In summary, anchored decomposition and separable decomposition share a common goal of simplifying high-dimensional functions by focusing on lower-dimensional interactions. We note that each separable component $\Psi_{l}^{j}$ can be interpreted as a localized aggregation of the anchored decomposition terms with anchor $c=0$. Specifically, $\Psi_{l}^{j}$ corresponds to the sum of all functions $f_{u,0}$ for which the index set $u$ contains $j$ and any $i\in u$ satisfies $i\geq j$ and $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq l$. This yields the identity
$\Psi_{l}^{j}(z_{j,l})=\sum_{\begin{subarray}{c}u\subseteq\{j,j+1,\dots,s\}\\ j\in u,\;\max_{i\in u}\mathrm{dist}_{\mathcal{G}}(i,j)\leq l\end{subarray}}f_{u,0}(z_{u}),$
which makes explicit the structural relationship between our separable representation (4) and the anchored decomposition. While anchored decomposition systematically includes all terms based on variable subsets, separable decomposition selectively neglects terms involving distant variables, leveraging graph-based distances. This distinction allows separable decomposition to achieve greater computational efficiency by emphasizing localized interactions, making it particularly suited for functions with decaying sensitivity over distance.
@@ 4 Sensitivity decay in linear-quadratic optimal control
Based on Setting 2.1, we define the state $x\in\mathbb{R}^{n}$ as being composed of $s\leq n$ substates $x=(z_{1},\dots,z_{s})$, where each substate $z_{j}$ may have an associated control $u_{j}\in\mathbb{R}^{m_{j}}$ for $j\in\{1,\dots,s\}$. Whenever a substate does not have an associated control, we set $m_{j}=0$. The collection of all controls is denoted by $u=(u_{1},\dots,u_{s})\in\mathbb{R}^{m}$, with $m=\sum_{j=1}^{s}m_{j}$. We denote by $\mathcal{U}_{d}=\ell^{\infty}(\mathbb{N},\mathbb{R}^{m})$ the space of all bounded control sequences in the discrete‐time case, and by $\mathcal{U}_{c}=L^{\infty}([0,\infty),\mathbb{R}^{m})$ the space of all essentially bounded control sequences in the continuous‐time case.
Given matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $Q\in\mathbb{R}^{n\times n}$, $Q\succeq 0_{n}$, and $R\in\mathbb{R}^{m\times m}$, $R\succ 0_{m}$, we consider an infinite-horizon optimal control problem formulated either in discrete time
$\begin{split}\min_{u\in\mathcal{U}_{d}}\;&J(x_{0},u)=\sum_{k=0}^{\infty}\ell(x(k),u(k)),\quad\text{s.t.}\;x(k+1)=f(x(k),u(k)),\;x(0)=x_{0},\end{split}$ (12)
@@
$\begin{split}\min_{u\in\mathcal{U}_{c}}\;&J(x_{0},u)=\int_{0}^{\infty}\ell(x(t),u(t))\,dt\,\quad\text{s.t.}\;\dot{x}(t)=f(x(t),u(t)),\;x(0)=x_{0},\end{split}$ (13)
@@
where $\mathcal{U}=\mathcal{U}_{d}$ in the discrete-time case and $\mathcal{U}=\mathcal{U}_{c}$ in the continuous-time case. In the LQR framework, it is well-known that the optimal value function $V$ is quadratic, taking the form $V(x)=x^{\top}Px$, where $P$ satisfies the Discrete Algebraic Riccati Equation for discrete-time systems, and the Continuous Algebraic Riccati Equation for continuous-time systems. As described in Setting 2.1, the graph $\mathcal{G}$ associated with such a problem is determined by the subsystems and their mutual influences. For the set of vertices $\{1,\dots,s\}$, the graph $\mathcal{G}$ has an edge $(j,i)$ if
@@
As highlighted in Corollary 3.1, the decaying sensitivity of the optimal value function is associated with the SED property of the Riccati solution. In the first subsection, we address the discrete-time setting, where we revisit results previously derived in [33]. In the second subsection, we transition to the continuous-time setting and, under appropriate assumptions on the problem data, demonstrate an off-diagonal decay property of the Riccati solution. We introduce some notions that will be used in the following.
@@
The pair $(A,B)$ is called stabilizable if there exists a feedback matrix $K\in\mathbb{R}^{m\times n}$ such that ABKA-BK is stable.
@@ 4.1 Discrete-Time Linear Quadratic Regulator
    $(A,B)$ is uniformly stabilizable, i.e., Definition 4.1 holds with $\|K\|\leq L$ and C=ABKC=A-BK satisfies (15),
@@
$\|K[i,j]\|\leq C_{K}\rho_{K}^{\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)},$
@@
Note that Proposition 4.1 is stated for the case $\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)=|i-j|$, but its proof extends naturally to any graph where $|\{j\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq 1\}|$ remains bounded independently of the dimension for all $i\in\{1,\dots,s\}$. For example, if $B=Q=R=I_{n}$ and $A$ is a $p$-banded matrix for some $p\in\mathbb{N}$, it follows that $|\{j\mid\operatorname{\operatorname{dist_{\mathcal{G}}}}(i,j)\leq 1\}|\leq 2p+1$ for any dimension and any node $i$.
@@ 4.2 Continuous-Time Linear Quadratic Regulator
First, observe that $A^{2}+\frac{c}{\gamma}I_{n}\succ 0$, ensuring that the matrix square root in the expression for $P$ is well defined. Furthermore, since $\sqrt{\lambda^{2}+\frac{c}{\gamma}}>|\lambda|$ and $\gamma>0$, the matrix in (18) is positive definite. Substituting expression (18) into the ARE (17) confirms that the expression for $P$ satisfies the equation. ∎
We now introduce the concept of functions of diagonalizable matrices, which will be instrumental in the subsequent analysis. For further details and alternative definitions, the reader is referred to [16].
Definition 4.5 (Function of a diagonalizable matrix).
Let $A\in\mathbb{R}^{n\times n}$ be a diagonalizable matrix. Then there exists an invertible matrix $X\in\mathbb{C}^{n\times n}$ and a diagonal matrix $\Lambda=\operatorname{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{n})\in\mathbb{C}^{n\times n},$ containing the eigenvalues of $A$, such that $A=X\Lambda X^{-1}.$ Let $f(z)$ be a scalar-valued function (real or complex) that is defined at each eigenvalue $\lambda_{i}$ of $A$. The function of the matrix $A$, denoted $f(A)$, is then defined by
$f(A)=Xf(\Lambda)X^{-1},$
where
$f(\Lambda)=\operatorname{diag}(f(\lambda_{1}),f(\lambda_{2}),\dots,f(\lambda_{n})).$
Then, the expression (18) can be reformulated as $P=f(A)$, where
@@
remains banded with the same bandwidth, and its spectrum is contained in [1,1][-1,1]. Moreover, $)f(A)=f(\psi^{-1}(\widetilde{A}$)). Accordingly, we focus on bounding the off-diagonal entries of the matrix function F(A~)F($\widetilde${A}), where for $x\in\mathbb{C}$
@@
and $\widetilde{A}$ is banded with spectrum in [1,1][-1,1]. The following useful result is adapted from [5, Theorem 2.2].
@@
Let $g:\mathbb{C}\rightarrow\mathbb{C}$ be a function analytical in the interior of an ellipse $\mathcal{E}_{\chi}$ with foci at $\pm 1$ and sum of the half axes $\chi>1$, and continuous on $\mathcal{E}_{\chi}$. Additionally, let $A\in\mathbb{C}^{n\times n}$ be symmetric, $p$-banded, and such that [1,1][-1,1] contains its spectrum, and let $\rho:=\chi^{-\frac{1}{p}}\in(0,1)$. Then, for all pairs of indices $(i,j)$ such that $i\neq j$, we have that
@@
and $\mathcal{E}_{\chi}$ is the ellipse with foci at $\pm 1$ and sum of the half axes $\chi>1$.
@@
$x_{1/2}=-\frac{b+a}{b-a}\pm i\,2\frac{\sqrt{c}}{\sqrt{\gamma}(b-a)}.$
Thus, we consider the ellipse $\mathcal{E}_{\chi}$ with foci in 1-1 and $1$ passing through the points $x_{1/2}$. The sum of the half axes of $\mathcal{E}_{\chi}$ is given by
@@ 5 Efficient Neural Network Representation for Separable Structures
In this section, we examine how neural networks can efficiently represent functions exhibiting decaying sensitivity. While neural networks are universal approximators, they typically require an exponential growth in neurons and parameters as input dimensionality increases. However, we demonstrate that for functions $V$ possessing a separable approximation, this curse of dimensionality can be mitigated. Under appropriate conditions, such functions can be represented with neural networks requiring a number of parameters which only grows polynomially. This result is consistent with existing literature showing that certain structural properties enable polynomial parameter scaling with dimension. For example, a similar result holds for compositional functions (cf. [17]) and in particular, for the sub-class of separable functions. The approximation of separable functions with neural networks has been discussed in [13, 32] in the context of (control) Lyapunov functions. While the approach and the neural network construction in the following theorem is similar, there are two major differences compared to these works. The first one is that [13, 32] considers strictly separable functions, meaning that each component $x_{i}$ of the state $x$ appears in exactly one subsystem. This is not the case in the framework of this paper, since we achieve separability by considering overlapping graph neighborhoods, whence we cannot directly apply the results in [13, 32] towards our setting. On the other hand, in [13, 32] the authors used a linear transformation as the first hidden layer for the computation of suitable substates to achieve separability. This is not necessary in our case, since for any given length $l$ of the graph-neighborhoods, the vectors $z_{j,l}$ are uniquely determined by the graph $\mathcal{G}$.
@@
respectively, where $C>0$ and $d>0$ are independent of $s$ and $n$, but are determined by $\varepsilon$, $R$, $\widetilde{C}$, $\delta$, and $d_{j,l}$, where $\widetilde{C}$ and $\delta$ are defined as in Theorem 3.4, and $d_{j,l}$ denotes the dimension of the subvectors $z_{j,l}$ introduced in Subsection 3.2, with $l$ itself depending on $\varepsilon$, $\widetilde{C}$, and $\delta$.
@@
Let $l\coloneqq\left\lceil{\ln\left(\epsilon{\widetilde{C}}^{-1}\right)\ln^{-1}(\delta)-1}\right\rceil\in\mathbb{N}$, where $\left\lceil{\cdot}\right\rceil$ stands for the ceiling function. Let $\Psi_{l}$ be as defined in (4) and (5). Then, by Theorem 3.4, we obtain that for all $x\in\Omega$
@@
Thus, it is sufficient to show the existence of a neural network that approximates $\Psi_{l}$ on $\Omega$ with error bounded by $\epsilon$. We begin by constructing subnetworks for approximating $\Psi_{l}^{j}$, $1\leq j\leq s$. Recall that $\Psi_{l}^{j}$ depends on $z_{j,l}\in\mathbb{R}^{d_{j,l}}$. Since $\sigma$ is non-polynomial, by Theorem 2.1 in [23] and its extension to cubes with length $R$ in [17], there exists a constant $\mu_{d_{j,l}}$ such that any neural network $W_{j}(\cdot;\theta_{j})\colon\mathbb{R}^{d_{j,l}}\to\mathbb{R}$ consisting of one hidden layer with $M\in\mathbb{N}$ neurons and activation function $\sigma$ satisfies
$\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\leq\mu_{d_{j},l}\,\widetilde{R}\,{M}^{-\frac{1}{d_{j,l}}}\|\Psi_{l}^{j}\|_{C^{1}([-R,R]^{d_{j,l}})},$
where $\widetilde{R}\coloneqq\max\{1,R\}$. Since, by assumption, $V$ has a $C^{1}$-norm bounded by $C_{V}$ on $R,R]^{n}$, it follows from (5) that $\Psi_{l}^{j}$ has $C^{1}$-norm bounded by $2C_{V}$ on $R,R]^{d_{j,l}}$, leading to
$\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\leq 2C_{V}\mu_{d_{j},l}\,\widetilde{R}\,{M}^{-\frac{1}{d_{j,l}}}.$ (22)
@@
$\begin{split}&\inf_{\theta}|W(x;\theta)-\Psi_{l}(x)|_{\infty,\Omega}\leq\sum_{j=1}^{s}\inf_{\theta_{j}}|W_{j}(z_{j,l};\theta_{j})-\Psi_{l}^{j}(z_{j,l})|_{\infty,[-R,R]^{d_{j,l}}}\\ \leq&2sC_{V}\mu_{d}\,\widetilde{R}\,{M}^{-\frac{1}{d}},\end{split}$
where $d\coloneqq\max_{j}d_{j,l}$. Solving for $M$ yields $M\geq s^{d}(2C_{V}\mu_{d}\widetilde{R})^{d}\frac{1}{\varepsilon^{d}}$. To establish the existence of a parameter $\theta$ that achieves the desired error bound $\varepsilon$, we choose
$M=s^{d}\left\lceil 2C_{V}\mu_{d}\widetilde{R}{\varepsilon}^{-1}\right\rceil^{d}=:s^{d}C^{d}$
@@
Theorem 5.1 states that, under the assumption of exponentially decaying sensitivity, neural networks can approximate $V$ with a number of neurons and parameters that grows polynomially with respect to the number of subsystems $s$, where the exponent is $d$. The maximum dimension of a graph-neighborhood $d$ depends on $l$, which is determined by $\varepsilon,\widetilde{C}$, and $\delta$, but independent of $s$ and $n$. Thus, for a family of problems with increasing dimension and uniform decaying sensitivity, where, for any fixed $l$, the maximum neighborhood dimension $d$ remains constant, Theorem 5.1 describes a situation where the curse of dimensionality is avoided. This applies, for example, to a sequential graph with one-dimensional subsystems, where $d=l+1$ for any number of subsystems. However, it is important to note that Theorem 5.1 does not address the required growth of the network in relation to the accuracy $\varepsilon$.